# Accelerated decomposition of bistochastic kernel matrices by low rank approximation

### v1

@@ 3.1 Dilution
We use the low rank approximation (1) to design an accelerated algorithm of reduced overall cost. To that end we begin by computing the EVD of $\bm{F}^{\top}\bm{F}\in\mathbb{R}^{r\times r}$
@@
$\tilde{\bm{P}}=\bm{G}\bm{G}^{\top}=(\bm{Q}_{1}\bm{U}_{1})\bm{\Sigma}_{1}^{2}(\bm{Q}_{1}\bm{U}_{1})^{\top}=\bm{\Phi}\bm{\Lambda}\bm{\Phi}^{\top}$ (3)
@@ 3.3 Discussion
On the other hand, the dilution method has the advantage of incorporating at least partial information about the whole dataset $X_{N}$. This is because the partial Cholesky factorization of $\bm{K}$ involves projecting the matrix onto the linear span of its $r$ selected columns. Thinking of each column $i\in\{0,\ldots,N-1\}$ as corresponding to state $x_{i}\in X_{N}$, every column $i$ sampled by the pivoted partial Cholesky algorithm contains the pairwise distances $k(x_{i},x_{j})$ between state $x_{i}$ and all other states $x_{j}\in X_{N}$, $j\in\{0,\ldots,N-1\}$. On the contrary, the subsampling method does not utilize any information about the states rejected during the sampling process. In Section 4.2 we demonstrate empirically two ways in which this difference between the two methods may manifest itself in the obtained eigenvalues.
@@ 4 Application
$\partial_{t}u=-u\partial_{s}u-\partial_{s}^{2}u-\partial_{s}^{4}u\qquad t\geq 0,\quad s\in S$ (4)
@@
In addition to its rich dynamics and well understood bifurcation diagram, the KS problem (4) has additional desirable properties that make it an excellent testbed for pattern extraction methods. First, its generated solutions satisfy symmetries acting on the spatial domain; more specifically, the KS problem (4) is equivariant under translations $u(t,x)\mapsto u(t,x+y)$ for all $y\in\mathbb{R}$ and reflection $u(t,x)\mapsto-u(t,-x)$. Second, it has a global attractor of finite dimension; namely, a finite dimensional subset $A\subset X$ which is forward invariant under the dynamics and attracts all initial conditions $u(0,\cdot)\in X\subset H_{s}$ [43, 40].
The dynamics generated by the KS problem (4) is given by the flow map $\Phi^{t}:X\to X$ with continuous time variable $t\geq 0$, state space $X$ and invariant probability measure $\mu$ with compact support $\operatorname{supp}\mu\subseteq A$. In what follows we also employ the discrete time flow map $\Phi^{n\Delta t}:X\to X$ with sampling timestep $\Delta t\geq 0$ and $n\in\mathbb{N}$. The real valued observables of the dynamics are members of the Hilbert space of functions $H_{X}=L^{2}(X,\mu;\mathbb{R})$.
@@ 4.1 Spatiotemporal pattern extraction
$Kf(\omega)=\int_{\Omega}\kappa(\omega,\omega^{\prime})f(\omega^{\prime})d\sigma(\omega^{\prime})$ (5)
@@
$k(W_{Q}(\omega),W_{Q}(\omega^{\prime})=\exp\bigl(-\frac{1}{\epsilon Q}\lVert W_{Q}(\omega)-W_{Q}(\omega^{\prime})\rVert_{2}^{2}\bigr)$ (6)
where $\epsilon>0$ is a tunable bandwidth parameter and $\lVert\cdot\rVert_{Q}$ denotes the 2-norm in $\mathbb{R}^{Q}$. For our numerical experiments (Section 4.2) we are going to use the bistochastic normalized versions of kernel (6) and associated kernel integral operator (5).
Figure 1: (Left) Space-time heatmap of the true state data obtained by integrating the KS problem (4) for $500$ time units using the parameter values given in Section 4.2. (Right) Space-time heatmap of the same state data as in the left plot with black dots indicating the states sampled by the pivoted partial Cholesky algorithm for rank parameter $r=2048$.
Thanks to its acting by composition with the delay embedding map $W_{Q}$, kernel function $\kappa$ factors the product state space $\Omega$ into equivalence classes consisting of states with identical dynamical behavior under $Q$ delays. As shown in [26], this implies that the functions in the range $\operatorname{ran}K$ of integral operator (5) are invariant under the actions of the symmetry group of the KS problem (4). To make this property precise, we consider the group of symmetries $G$ with continuous left action on the spatial domain $\Gamma_{S,g}\colon S\to S$ for $g\in G$. For the KS problem (4), the actions $\Gamma_{S,g}$ represent spatial translations and reflection. Every induced action $\Gamma_{X,g}\colon X\to X$, $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$ represents a dynamical symmetry of the dynamics generated by (4), meaning that the dynamics $\Phi^{t}$ satisfies the equivariance property
@@
$f\circ\Gamma_{\Omega,g}=f$ (7)
for all $g\in G$ with induced action $\Gamma_{\Omega,g}\colon\Omega\to\Omega$, $\Gamma_{\Omega,g}=\Gamma_{X,g}\otimes\Gamma_{S,g}$. Thanks to the invariance property (7), every eigenfunction of the kernel integral operator (5) is invariant under the actions of $G$, meaning that each such function can generally represent a more complex spatiotemporal pattern than when symmetry invariance is not ensured [26].
@@ 4.2 Numerical experiments
For our numerical experiments we consider the KS problem (4) for the domain length value $L=22$ which corresponds to chaotic dynamics. We perform a Fourier spatial discretization using $M=64$ Fourier modes and 3/2 dealiasing of the pseudospectral treatment of the quadratic nonlinearity [9]. For the temporal discretization we employ the exponential time differencing 4-stage Runge-Kutta (ETDRK4) method introduced in [15] and further developed in [31], using the timestep value $\Delta t=0.25$. Our initial condition is formed by setting the leading four Fourier coefficients to $0.6$ and the rest to zero. After integrating for 10 00010\,000 timesteps (2 5002\,500 time units) we collect one sample every four timesteps (one time unit) for a total of $563$ samples. A space-time plot of the integrated solution is shown in the left panel of Figure 1.
@@
Using $Q=64$ delays, our training dataset consists of $N=500$ samples in delay embedded form, bringing the total number of product state samples to NM=32 500NM=32\,500. To employ the dilution and subsampling algorithms developed in the previous section we must first calibrate the bandwidth parameter $\epsilon$ of kernel (6). We perform the bandwidth calibration in two stages. First, we use the median rule studied in [24] and the full training dataset to get a first approximation of parameter $\epsilon$. Second, using the obtained $\epsilon$ value we apply the pivoted partial Cholesky algorithm developed in [11] to sample $8192$ states from our training dataset. Relying only on the sampled subset of training data, we use the scaling-based algorithm developed in [14] to refine our approximation of bandwidth $\epsilon$. The bandwidth value selected by this two-stage algorithm is $\epsilon=15$. After experimenting with similar bandwidth values, the numerical results presented below are obtained using the adjusted value $\epsilon=50$.
@@
Next we employ the dilution and subsampling algorithms to approximate the EVD of the kernel matrix corresponding to the bistochastic normalized version of kernel (6). For that we perform a partial Cholesky factorization with rank parameter $r=2048$, yielding a trace norm error of $7.15\%$. The right panel of Figure 1 shows the product states sampled by the pivoted partial Cholesky algorithm for the used rank parameter.
@@
Returning to the eigenvalues comparison in Figure 5, one noticeable difference between the two sets of eigenvalues is their rate of decay as the eigenvalue index is growing. In particular, the subsampling eigenvalues have a tail that decays more quickly than the dilution eigenvalues. This is reasonable considering that the subsampling method uses only a subset of the original samples and does not contain any information about the rejected samples. On the contrary, the dilution method employs at least partial information about all training samples (Section 3.3). This seems to translate to a wider eigenvalue spectrum for the dilution eigenvalues; namely, a slower rate of decay of the magnitude of the eigenvalues which is closer to the true decay rate.
In addition, it appears that the subsampling method overestimates the magnitude of the eigenvalues for moderate indices, compared to the dilution and true results. Combined with the faster decay rate for large indices, this can be interpreted as the subsampling method placing a larger relative importance to the eigenfunctions corresponding to low and moderate indices. In this line of thinking, the relative “importance” of an eigenfunction is defined as the ratio of its eigenvalue over the sum of all eigenvalues, qualitatively similar to the “energy ratio” used in proper orthogonal decomposition [2]. This can again be thought of as a consequence of the fact that the subsampling method does not incorporate any information about the states rejected in the sampling process.
@@ 4.3 Discussion
Based on our comparison of the eigenvalues and eigenfunctions obtained by each of the dilution and subsampling methods, we conclude that the dilution method can be expected to yield a more accurate approximation for a given approximation rank. This is reflected not only on the closer agreement of the dilution eigenfunctions with the true ones, but also on the more accurate approximation of the true eigenvalue spectrum. This property seems to be a result of the fact that the dilution method incorporates at least partial information about all state samples in the training dataset.
@@
The choice of kernel function used in defining the kernel integral operator (5) plays an important role in determining the rank of its range, and ultimately the maximum rank that can be obtained by the linear span of the numerically computed eigenfunctions. Our use of a gaussian kernel means that the rank of the integral operator is infinite, regardless of the rank of the observation map used to generate the employed state data [41, 26]. On the contrary, the state correlation kernels traditionally used in POD do not share that property, meaning that the rank of the associated integral operator is bounded by the rank of the observation map. Although the rank obtained by POD is sufficient to represent the state data to any given level of accuracy, it might not be sufficient to represent arbitrary observables. As a result, our choice of the gaussian kernel (6) is motivated by applications where representing observables based on the computed eigenfunctions is important [22, 23, 46].
Finally, for our numerical results we used the gaussian kernel (6) with a fixed bandwidth parameter $\epsilon$. The methods considered in this work can be applied without change to kernels of variable bandwidth, where the employed bandwidth depends on the arguments of the kernel function [7, 25].

### v2

@@ 1.2 Partial Cholesky factorization
For an $N\times N$ positive semidefinite matrix $\bm{A}$ and a rank parameter $r^{\prime}\in\mathbb{N}$ we denote by $\bm{A}]_{r^{\prime}}\in\mathbb{R}^{N\times N}$ a best rank-$r^{\prime}$ approximation of $\bm{A}$, which is generally not unique. As before, we denote by $\tilde{\bm{A}}\in\mathbb{R}^{N\times N}$ the rank-$r$ approximation produced by the randomly pivoted partial Cholesky algorithm with the sampling strategy of [11], for a rank parameter $r\geq r^{\prime}$. For fixed parameters $r^{\prime}$ and $\epsilon>0$ it is shown in [11] that the approximation $\tilde{\bm{A}}$ satisfies the error bound
$\mathbb{E}\,\frac{\operatorname{tr}{(\bm{A}-\tilde{\bm{A}})}}{\operatorname{tr}{\bm{A}}}\leq(1+\epsilon)\,\eta\qquad\eta=\operatorname{tr}{(\bm{A}-[\bm{A}]_{r^{\prime}})}/\operatorname{tr}{\bm{A}}$ (1.1)
for every $r$ such that
$r\geq\frac{r^{\prime}}{\epsilon}+r^{\prime}\log\frac{1}{\epsilon\eta}$
where the average is taken over all randomly sampled columns. Namely, when the number of columns $r$ of the partial Cholesky factorization satisfies the above inequality, then the relative trace norm error of the approximation is at most $\epsilon$ times greater than the best relative error for an approximation of lower rank $r^{\prime}\leq r$. We refer the reader to [11] for a precise statement of the proven error bounds and comparison to other sampling strategies. We will use the above approximation error bound to offer an explanation for some of the numerical results presented in Section 4.2.
@@ 3.1 Dilution
We use the low rank approximation (3.1) to design an accelerated algorithm of reduced overall cost. To that end we begin by computing the EVD of $\bm{F}^{\top}\bm{F}\in\mathbb{R}^{r\times r}$
@@
$\tilde{\bm{P}}=\bm{G}\bm{G}^{\top}=(\bm{Q}_{1}\bm{U}_{1})\bm{\Sigma}_{1}^{2}(\bm{Q}_{1}\bm{U}_{1})^{\top}=\bm{\Phi}\bm{\Lambda}\bm{\Phi}^{\top}$ (3.3)
@@ 3.3 Discussion
On the other hand, the dilution method has the advantage of incorporating at least partial information about the whole dataset $X_{N}$. This is because the partial Cholesky factorization of $\bm{K}$ involves projecting the matrix onto the linear span of its $r$ selected columns. Thinking of each column $i\in\{0,\ldots,N-1\}$ as corresponding to state $x_{i}\in X_{N}$, every column $i$ sampled by the pivoted partial Cholesky algorithm contains the pairwise distances $k(x_{i},x_{j})$ between state $x_{i}$ and all other states $x_{j}\in X_{N}$, $j\in\{0,\ldots,N-1\}$. For the choice of sampling strategy introduced in [11], this is also reflected in the error bound (1.1) satisfied by the dilution method’s approximation $\tilde{\bm{K}}$. On the contrary, the subsampling method does not utilize any information about the states rejected during the sampling process, and employs a kernel approximation $\tilde{\bm{K}}$ that is not guaranteed to satisfy the appropriate error bound (1.1). In Section 4.2 we demonstrate empirically two ways in which these differences between the two methods may manifest themselves in the obtained eigenvalues.
@@ 4 Application
$\partial_{t}u=-u\partial_{s}u-\partial_{s}^{2}u-\partial_{s}^{4}u\qquad t\geq 0,\quad s\in S$ (4.1)
@@
In addition to its rich dynamics and well understood bifurcation diagram, the KS problem (4.1) has additional desirable properties that make it an excellent testbed for pattern extraction methods. First, its generated solutions satisfy symmetries acting on the spatial domain; more specifically, the KS problem (4.1) is equivariant under translations $u(t,x)\mapsto u(t,x+y)$ for all $y\in\mathbb{R}$ and reflection $u(t,x)\mapsto-u(t,-x)$. Second, it has a global attractor of finite dimension; namely, a finite dimensional subset $A\subset X$ which is forward invariant under the dynamics and attracts all initial conditions $u(0,\cdot)\in X\subset H_{s}$ [43, 40].
The dynamics generated by the KS problem (4.1) is given by the flow map $\Phi^{t}:X\to X$ with continuous time variable $t\geq 0$, state space $X$ and invariant probability measure $\mu$ with compact support $\operatorname{supp}\mu\subseteq A$. In what follows we also employ the discrete time flow map $\Phi^{n\Delta t}:X\to X$ with sampling timestep $\Delta t\geq 0$ and $n\in\mathbb{N}$. The real valued observables of the dynamics are members of the Hilbert space of functions $H_{X}=L^{2}(X,\mu;\mathbb{R})$.
@@ 4.1 Spatiotemporal pattern extraction
$Kf(\omega)=\int_{\Omega}\kappa(\omega,\omega^{\prime})f(\omega^{\prime})d\sigma(\omega^{\prime})$ (4.2)
@@
$k(W_{Q}(\omega),W_{Q}(\omega^{\prime})=\exp\bigl(-\frac{1}{\epsilon Q}\lVert W_{Q}(\omega)-W_{Q}(\omega^{\prime})\rVert_{2}^{2}\bigr)$ (4.3)
where $\epsilon>0$ is a tunable bandwidth parameter and $\lVert\cdot\rVert_{Q}$ denotes the 2-norm in $\mathbb{R}^{Q}$. For our numerical experiments (Section 4.2) we are going to use the bistochastic normalized versions of kernel (4.3) and associated kernel integral operator (4.2).
Figure 1: (Left) Space-time heatmap of the true state data obtained by integrating the KS problem (4.1) for $500$ time units using the parameter values given in Section 4.2. (Right) Space-time heatmap of the same state data as in the left plot with black dots indicating the states sampled by the pivoted partial Cholesky algorithm for rank parameter $r=2048$.
Thanks to its acting by composition with the delay embedding map $W_{Q}$, kernel function $\kappa$ factors the product state space $\Omega$ into equivalence classes consisting of states with identical dynamical behavior under $Q$ delays. As shown in [26], this implies that the functions in the range $\operatorname{ran}K$ of integral operator (4.2) are invariant under the actions of the symmetry group of the KS problem (4.1). To make this property precise, we consider the group of symmetries $G$ with continuous left action on the spatial domain $\Gamma_{S,g}\colon S\to S$ for $g\in G$. For the KS problem (4.1), the actions $\Gamma_{S,g}$ represent spatial translations and reflection. Every induced action $\Gamma_{X,g}\colon X\to X$, $\Gamma_{X,g}(x)=x\circ\Gamma_{S,g}^{-1}$ represents a dynamical symmetry of the dynamics generated by (4.1), meaning that the dynamics $\Phi^{t}$ satisfies the equivariance property
@@
$f\circ\Gamma_{\Omega,g}=f$ (4.4)
for all $g\in G$ with induced action $\Gamma_{\Omega,g}\colon\Omega\to\Omega$, $\Gamma_{\Omega,g}=\Gamma_{X,g}\otimes\Gamma_{S,g}$. Thanks to the invariance property (4.4), every eigenfunction of the kernel integral operator (4.2) is invariant under the actions of $G$ on $\Omega$, meaning that each such function can generally represent a more complex spatiotemporal pattern than when symmetry invariance is not ensured [26].
@@ 4.2 Numerical experiments
For our numerical experiments we consider the KS problem (4.1) for the domain length value $L=22$ which corresponds to chaotic dynamics. We perform a Fourier spatial discretization using $M=64$ Fourier modes and 3/2 dealiasing of the pseudospectral treatment of the quadratic nonlinearity [9]. For the temporal discretization we employ the exponential time differencing 4-stage Runge-Kutta (ETDRK4) method introduced in [15] and further developed in [31], using the timestep value $\Delta t=0.25$. Our initial condition is formed by setting the leading four Fourier coefficients to $0.6$ and the rest to zero. After integrating for 10 00010\,000 timesteps (2 5002\,500 time units) we collect one sample every four timesteps (one time unit) for a total of $563$ samples. A space-time plot of the integrated solution is shown in the left panel of Figure 1.
@@
Using $Q=64$ delays, our training dataset consists of $N=500$ samples in delay embedded form, bringing the total number of product state samples to NM=32 500NM=32\,500. To employ the dilution and subsampling algorithms developed in the previous section we must first calibrate the bandwidth parameter $\epsilon$ of kernel (4.3). We perform the bandwidth calibration in two stages. First, we use the median rule studied in [24] and the full training dataset to get a first approximation of parameter $\epsilon$. Second, using the obtained $\epsilon$ value we apply the pivoted partial Cholesky algorithm developed in [11] to sample $8192$ states from our training dataset. Relying only on the sampled subset of training data, we use the scaling-based algorithm developed in [14] to refine our approximation of bandwidth $\epsilon$. The bandwidth value selected by this two-stage algorithm is $\epsilon=15$. After experimenting with similar bandwidth values, the numerical results presented below are obtained using the adjusted value $\epsilon=50$.
@@
Next we employ the dilution and subsampling algorithms to approximate the EVD of the kernel matrix corresponding to the bistochastic normalized version of kernel (4.3). For that we perform a partial Cholesky factorization with rank parameter $r=2048$, yielding a trace norm error of $7.15\%$. The right panel of Figure 1 shows the product states sampled by the pivoted partial Cholesky algorithm for the used rank parameter.
@@
Returning to the eigenvalues comparison in Figure 5, one noticeable difference between the two sets of eigenvalues is their rate of decay as the eigenvalue index is growing. In particular, the subsampling eigenvalues have a tail that decays more quickly than the dilution eigenvalues. This is reasonable considering that the subsampling method uses only a subset of the original samples and does not contain any information about the rejected samples. On the contrary, the dilution method employs at least partial information about all training samples (Section 3.3). In this case, this translates to a wider eigenvalue spectrum for the dilution eigenvalues; namely, a slower rate of decay of the magnitude of the eigenvalues which is closer to the true decay rate.
In addition, it appears that the subsampling method overestimates the magnitude of the eigenvalues for moderate indices, compared to the dilution and true results. Combined with the faster decay rate for large indices, this can be interpreted as the subsampling method placing a larger relative importance to the eigenfunctions corresponding to low and moderate indices. In this line of thinking, the relative “importance” of an eigenfunction is defined as the ratio of its eigenvalue over the sum of all eigenvalues, qualitatively similar to the “energy ratio” used in proper orthogonal decomposition [2].
The above empirical observations are backed by the fact that the rank-$r$ kernel approximation $\tilde{\bm{K}}$ employed by the dilution method satisfies the error bound (1.1) for appropriate choices of parameters $r^{\prime}\leq r$ and $\epsilon$, which is generally not true for the subsampling method. This implies that the dilution method’s $\tilde{\bm{K}}$ is close to a spectrally optimal approximation of $\bm{K}$ of rank $r^{\prime}\leq r$, which is reflected in the very close agreement between the dilution and true eigenvalues and eigenfunctions for low to moderate indices.
@@ 4.3 Discussion
Based on our comparison of the eigenvalues and eigenfunctions obtained by each of the dilution and subsampling methods, we conclude that the dilution method can be expected to yield a more accurate approximation for a given approximation rank. This is reflected in the closer agreement of the dilution eigenvalues and eigenfunctions with the true ones, as well as in the approximation error bound (1.1) satisfied by the low rank approximation $\tilde{\bm{K}}$ employed by the dilution method.
@@
The choice of kernel function used in defining the kernel integral operator (4.2) plays an important role in determining the rank of its range, and ultimately the maximum rank that can be obtained by the linear span of the numerically computed eigenfunctions. Our use of a gaussian kernel means that the rank of the integral operator is infinite, regardless of the rank of the observation map used to generate the employed state data [41, 26]. On the contrary, the state correlation kernels traditionally used in POD do not share that property, meaning that the rank of the associated integral operator is bounded by the rank of the observation map. Although the rank obtained by POD is sufficient to represent the state data to any given level of accuracy, it might not be sufficient to represent arbitrary observables. As a result, our choice of the gaussian kernel (4.3) is motivated by applications where representing observables based on the computed eigenfunctions is important [22, 23, 46].
Finally, for our numerical results we used the gaussian kernel (4.3) with a fixed bandwidth parameter $\epsilon$. The methods considered in this work can be applied without change to kernels of variable bandwidth, where the employed bandwidth depends on the arguments of the kernel function [7, 25].